Calculate the velocity of sound in air at 50oC, if the velocity of sound at 0oC is 330m/s when the pressure is doubled.


1
Expert's answer
2021-07-30T11:18:03-0400
"v(50)=330+0.606(50)=360.3\frac{m}{s}"